the area of the quadrilateral formed by the lines y=2x+3, y=0, x=4, x=6 is

Grade:12

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Given line is y=2x+3
if x=4,y=2(4)+3=11 so p1=(4,11)
if x=6,y=2(6)+3=15 so p2=(6,15)
also p3=(4,0)
and p4=(6,0)
Now area of the quadrilateral can be found by the formula
=1/2\begin{vmatrix}(4)(6)(4)(6) &\\(0)(0)(11)(15)& \end{vmatrix}
=1/2(0+66+60-66)
=30 sq.units
Thanks and Regards
Shaik aasif
askIITians faculty